OpenStudy (anonymous):

If f(b)=cos2B+2sinB, the numerical value of f (pi/6) is what?

OpenStudy (anonymous):

Simply plug pi/6 in B

OpenStudy (anonymous):

do i put my calculator in radian?

OpenStudy (anonymous):

Like this: \[\Large \cos \left(2 * \frac{\pi}{6}\right) + 2 \sin \left(\frac{\pi}{6}\right)\] And yes, put it in radian.

OpenStudy (anonymous):

i got 1 1/2 but thats not an answer choice

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

omg i was up to late x_X its 3/2

OpenStudy (anonymous):

Haha, yeah, it's the same.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

lol oh my yes, wait another one, this one confuses me beyond belief , sin(arccos 1/2)

OpenStudy (anonymous):

It's a little tough, of course. Do you have arccos in your calculator?

OpenStudy (anonymous):

is it radical 3/2?

OpenStudy (anonymous):

yea lol i think i just found the answer i was just studying too late last night and too stressed out to focus x.x

OpenStudy (anonymous):

First, calculate arccos(½) then calculate sin( Ans )

OpenStudy (anonymous):

arccos(½) = pi / 3 so sin(arccos 1/2) = sin( pi / 3 )

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

The answer is 0.866, by the way.

OpenStudy (anonymous):

uhmm no my bro said just put it together and i got .866 and radical3/2 is equal to it

OpenStudy (anonymous):

LOL yea i get it

OpenStudy (anonymous):

Glad you got it. Do you have another question?

OpenStudy (anonymous):

me too lol and yes! i will close this and open it!